Hello, I'm not so sure what the problem is trying to tell me. I know that I need to graph it, but I'm not sure how to start the equation or what the input and output should be. I really need some help. Thank you so much! I'll be so happy for some help to understand this situation!
The height of a golf ball in meters can be described by the equation h=-4.9t^2+22.7t, where t is the number of seconds after it was hit. Use a calculator or spreadsheet to make a table of inputs and outputs for this function. The variable t should start out at zero and increment by 0.2.

Answer by Boreal(15235)

t=======h
0=======0, substitute 0 in for t
0.2=====-4.9(0.04)+22.7(0.2)=4.344; the 0.04 is 0.2^2.
0.4=====-4.9*(0.4)^2+22.7(0.4)=8.296
Just continue along these lines. Be sure you watch for the negative sign in front of the first term and square the value before multiplying by -4.9.
Calculator can do this with lists, and a spreadsheet can do this by writing
=[-4.9(x)^2+(22.7)*x] and making a column with x=0,0.2,0.4.